# Spontaneous orbital selective Mott phase in the two band Hubbard model

### v1

@@ III.2 Bands and interorbital hybridization
Figure 2: Lower Hubbard bands at the high symmetry points for total electron density $n=1.72$ for (A) $\lambda=0.06t$ (B) $\lambda=0.2t$ (C) $\lambda=0.36t.$ These respectively corresponds to the blue circle, square and star from Fig. (1). The colors indicate the character orbital nature of the band. In the OSMP phase, one band is below $E_{F}$ as expected and has mostly distinct feature of a single orbital. Whereas in the orbitally uniform phase phase, both bands cross Fermi-energy and are completely hybridized. Inset shows the upper Hubbard bands.
@@
We present the lower Hubbard bands in the main panels of Fig.(2) for three different $\lambda$. In the OSMP phase, one of the bands flattens and remains below the Fermi energy $E_{F}$, while in the OUP phase, both bands cross $E_{F}$. The color of the lines shows the difference of $x$ and $y$ spectral function $\delta A=A_{xx}(\mathbf{k},\omega)-A_{yy}(\mathbf{k},\omega)$ thus characterizing the orbital character of the bands if any. The bands are weakly hybridized at the intersection points for $\lambda=0.06t$ in Fig. (2A). This shows that although the hybridization between the $x$ and $y$ orbitals vanishes at the Fermi level, it remains finite at lower energies. In Fig. (2B), the hybridization increases as interorbital hopping increases. Furthermore the Mott band approaches the Fermi-energy in Fig. (2B) Finally, when we transition to an orbital uniform phase, we lose the orbital character of the electrons as shown in Fig. (2C). We present the upper Hubbard bands as an inset of each panel, and they are separated from the lower Hubbard band by an energy gap of $\sim U$. Interestingly, the upper Hubbard bands remain entirely unhybridized in the OSMP phase.
@@
Figure 3: Orbitally resolved spectral function of the lower Hubbard bands for $\lambda=0.06t$ and $n=1.72$, corresponding to the blue circle in the phase diagram in Fig. (1A) displays the intraorbital spectral function $A_{\alpha\alpha}$ as a function of energy $\omega-\mu$ for different high symmetry $k$-points. The spectral for each $k$-points are shifted by $1.5t$ for clarity. (B) Shows the interorbital spectral function $A_{\alpha\beta}$ for the same parameter. Here the results for each $k$-points are shifted by $0.3t$ for clarity.
@@
Therefore, we reveal the characteristics of the bands in the OSMP and OUP phases. First, in the OSMP phase, one of the bands flattens and goes below the Fermi level. Secondly, in the OSMP phase, the two bands have weak hybridization and retain their orbital character. Thirdly, the hybridization occurs only around specific momentum points where the two bands overlap. However, such hybridization remains appreciable for any finite $\lambda$, indicating the absence of a complete breakdown of hybridization in the OSMP phase.
@@ III.3 Fermi surface
Figure 4: (A) Displays the spectral function $A(\mathbf{k},E_{F})$ at $\lambda=0.06t$ in the OSMP phase for $n=1.72$ (blue circle on Fig. (1A) ). (B) Depicts the same for $\lambda=0.2t$ (blue square on Fig. ( 1A)), in the metastable region of the OSMP. (C) Fermi surface of the orbital uniform phase at $\lambda=0.36t$ (blue star on Fig. (1A)). In the OSMP, only a single Fermi sheet is observable, whereas in the metallic phase, two sheets are visible. Furthermore, the Fermi surface area is smaller in the OSMP phase compared to the OUP. (D) Illustrates the evolution of the Fermi volume $\mathcal{A}_{s}$ as a function of interorbital hopping for different electron densities, showing a jump at the transition. In the OSMP phase, the Mott orbital does not contribute to the Fermi volume since one band remains filled.
In Fig. (4A), we illustrate the Fermi surface contour, for $\lambda=0.06t$ at $n=1.72$ in the OSMP phase. In the OSMP phase, a single Fermi sheet is evident since the Mott orbital remains fully occupied. With increasing interorbital hopping there is little change in the Fermi surface as shown in Fig. (4B). In contrast, both bands contribute to the Fermi surface in the metallic regime at $\lambda=0.36t$. Additionally, there is a significant increase in the size of the Fermi surface from the OSMP to the orbitally uniform phase state.
Moreover, we monitor the increase in the volume of the Fermi surface across the transition with interorbital hopping for various total electron densities in Fig. (4D). The volume exhibits a distinct jump at $\lambda_{c}$, which is more pronounced for lower $n$. From the evolution of the Fermi surface, the OSMP phase appears to be in a Kondo breakdown-like phase. However, we have shown in the previous section that hybridization of the two orbitals remains non-zero in the OSMP phase.
A characteristic feature of non-Fermi liquid behavior is the violation of the Luttinger theorem, which asserts that the volume of the Fermi surface must equal the electron density. This theorem [48] applies to any system characterized by a low-energy theory of a Fermi liquid. As the composite excitations in our system are non-fermionic, a violation of the Luttinger theorem is expected [49, 50]. Consequently, our calculations demonstrate that the Luttinger theorem is violated in both the OSMP and strongly correlated OUP phase.
Hence, we observe a small Fermi surface in the OSMP to a large Fermi surface in the OUP phase. As a result, we detect a violation of the Luttinger theorem in the orbital selective phase. Despite the sharp increase in the Fermi volume in the OUP phase, the Luttinger theorem continues to be violated. Thus, the traditional Fermi liquid picture is not applicable in the strongly correlated Hubbard model within this formalism.
@@ III.4 Density of states
Figure 5: (A) Presents the low-energy density of states at $n=1.72$ for several $\lambda$ ranging from 00 to $0.42t$ with a $0.02t$ step. Each $\lambda$ curves are shifted vertically by $0.05t$ for clarity.$\lambda<0.25t$ corresponds to the OSMP phase. (B), shows the orbital resolved density of states at $\lambda=0.16t$. Here, the $y$ orbital approaches the Mott character, indicated by the reduced DOS at Fermi energy, whereas the $x$ orbital remains metallic. The peak at the negative energy comes from the Mott orbital. A progressive flattening of such peak of the OSMP occurs in (A) with increasing interorbital hopping. (C) Shows the value of DOS at Fermi-energy, which manifest a jump around $\lambda=0.24t$, indicating the transition from OSMP to the orbitally uniform phase state. (D) Tracks the energy associated with the Mott peak $\omega_{\rm max}$ (black dots in A).
We now focus on the density of states (DOS). In Fig. (5A), we illustrate the low-energy density of states for $n=1.72$ across different $\lambda$ values. The curves for different $\lambda$ values are shifted vertically by $0.05t$ for clarity. Initially, at weak $\lambda$, the DOS features a prominent peak at negative energies (shown in black dots), with a finite DOS at the Fermi energy $E_{F}$. The presence of mobile electrons in one orbital maintains an ungapped DOS at $E_{F}$. As interorbital hopping increases, so does the DOS at Fermi energy. The transition from the OSMP to the OUP state is marked by a jump in the evolution of $D(E_{F})$, as shown in Fig. (5C).
In Fig. (5B), we depict the orbital-resolved DOS for ${\lambda=0.17}$. The Mott orbital $y$ displays a distinct peak at negative energy, accompanied by a small DOS at $E_{F}$. This peak signifies the energy required to extract electrons from the Mott orbital. We track the evolution of the Mott peak $\omega_{\rm max}$ as a function of interorbital hopping in Fig. (5D). The energy $\omega_{\rm max}$ shows a smooth reduction. As a result of the hybridization between heavy Mott orbital electrons and itinerant electrons, the peak flattens, making it challenging to discern the orbital character of electrons prior to the transition to OUP.
Therefore, the OSMP phase features a reduced DOS at Fermi energy. The transition from the OSMP to the orbitally uniform phase is marked by a sharp increase in the states at the Fermi energy at a critical interorbital hopping. Additionally, the OSMP phase displays a strong negative bias peak, which flattens, indicating the hybridization of Mott electrons with itinerant electrons. Notably, the flattening of the Mott peak with increasing interorbital hopping occurs more smoothly than the sharp transitions observed in other quantities during the phase transition. This indicates that hybridization among the orbitals ultimately eliminates the orbital differentiation.
@@ III.5 Quasiparticle residue
Figure 6: (A)) Quasiparticle weight for several density as a function of $\lambda$, the errorbars are standard deviation when averaging over Fermi momentum. We overlapped with points the value of the jumps observable on $N_{k}$. (B) orbital resolved quasiparticle weight. The Mott orbital has a vanishing quasiparticle residue for finite $\lambda$.
@@
We also report the typical shape of $N_{k}$ along the $\Gamma$ to $M$ high symmetry line in Fig. (8). By definition of the Green’s function in Eq. 79, when crossing the Fermi contour a discontinuity of the order of $Z$ is observed in momentum space. One drop is observed for the OSMP, and the momentum resolved density does not vanish at $M$ due to the presence of the fully filled Mott band. Two drops are seen in the OUP due to the two Fermi contours of the phase. In this case, $Z$ is defined as the sum of these two drops. We report the $Z$ measured this way as a function of $\lambda$ on Fig (6A) as dots. There is a good agreement between the two ways of computing $Z$. Since there is no Fermi surface at half-filling, and that $Z$ is a quantity defined at the Fermi energy, $Z=0$ at half-filling. As electron density $n$ is decreasing, $Z$ increases. Therefore, we can also interpret the quasi-particle weight as a measurement of how close the system is from Mott regime, where interaction affect the system the most.
Fig. (6B) displays the orbitally resolved quasi-particle weight. In the OSMP the Mott orbital has an almost (exactly at $\lambda=0$) vanishing quasiparticle weight due to the localized nature of the electron. Since the other orbital is further away from half-filling than the average doping its associated $Z$ increases from the average value.
@@ III.6 Phase diagram with Hubbard interactions
Figure 7: Here, we present the phase diagram as a function of intra-orbital interaction strength $U$ and $\lambda$ for $n=1.85$. The colored area shows the region where the OSMP phase is energetically favorable, whereas the OUP phase is favorable in the white region. The OSMP is a metastable state between the thick and dashed lines.
@@
Finally, In Fig. (7), we present the phase diagram when the on-site repulsion $U$ is varied for electron density $n=1.85$. The colored region indicates where the OSMP phase is more favorable than the uniform phase. On the other hand, the OUP phase is stable in the white areas. The OSMP is a metastable state between the thick and dashed lines.
@@
For lower $U$, there is a weak increment of critical interorbital hopping $\lambda_{c}$ where the OSMP phase becomes unfavorable. For large $U>15$ the same $\lambda_{c}$ reduces with increasing $U$. Thus, the critical $\lambda_{c}$ shows a nonmonotonic behavior with Hubbard interaction $U$.
@@
To qualitatively understand this result, we examine the destabilization of the OSMP phase. Initially, we assume double occupancy is forbidden due to strong repulsion. First, we focus on comprehending the decrease of $\lambda_{c}$ with increasing Hubbard interaction.
@@
In the OSMP phase, all the Mott orbital sites are singly occupied, whereas the metallic orbital has itinerant electrons and vacancies. When $\lambda\neq 0$, electrons can hop from the Mott to the metallic orbitals if the metallic site is vacant. A stable OSMP phase depends on refilling the holes in the Mott orbital without significant hybridization. Without recombination, holes will proliferate in the Mott orbital, leading to the breakdown of the orbital differentiation. The hole created in the Mott orbital is heavy and can be assumed to be approximately immobile.
@@
Two scenarios are possible once an electron hops to the metallic orbital: it either immediately hops back to the Mott orbital or transfers to a neighboring site within the same orbital. The stability of the OSMP phase depends on the ratio between $t$ and $\lambda$ for a fixed $U$ and electron density.
@@
In the $\lambda\ll t$ regime, electrons are more likely to hop between neighboring sites of the same orbitals rather than between orbitals. Since double occupancy is forbidden, recombination can only happen when any electron on the metallic orbital returns to the site at which Mott orbital has the hole and goes back to the Mott orbital. The mobility of electrons in the metallic orbitals is crucial in such a recombination pathway, which is relatively enhanced for lower $U$. Thus, the OSMP phase survives for a broader range of $\lambda$ when $U$ is lower.
@@
In the $\lambda\gg t$ regime, inter-orbital hopping between the $x$ and $y$ orbitals dominates. This leads to strong hybridization between the $x$ and $y$ orbitals, which destroys the OSMP phase.
@@
Finally, we focus on understanding why $\lambda_{c}$ decreases for $U<15$. In this regime, double occupancy contributes to destabilizing the OSMP ordering. As $U$ decreases, double occupancies become more likely, increasing hybridization between the Mott and itinerant orbitals. Consequently, a metastable OSMP phase can still be present at lower $U$ due to the competing processes that are favored when lowering $U$. However, the OSMP becomes energetically less favorable compared to the orbitally uniform phase.

### v2

@@ III.2 Bands and inter-layer hybridization
Figure 3: Lower Hubbard bands at the high symmetry points for total electron density $n=1.72$ for (A) $\lambda=0.06t$ (B) $\lambda=0.2t$ (C) $\lambda=0.36t.$ These respectively corresponds to the blue circle, square and star from Fig. (1). The colors indicate the layer character of the band at a particular momentum. In the LSMP phase, one band is below $E_{F}$ as expected and has mostly distinct feature of a single layer. Whereas in the uniform phase, both bands cross Fermi-energy and are completely hybridized. Inset shows the upper Hubbard bands.
@@
We present the lower Hubbard bands in the main panels of Fig.(3) for three different $\lambda$. In the LSMP phase, one of the bands flattens and remains below the Fermi energy $E_{F}$, while in the LUP phase, both bands cross $E_{F}$. The color of the lines shows the difference of $x$ and $y$ spectral function $\delta A=A_{xx}(\mathbf{k},\omega)-A_{yy}(\mathbf{k},\omega)$ thus characterizing the layer character of the bands if any. The bands are weakly hybridized at the intersection points for $\lambda=0.06t$ in Fig. (3A). This shows that although the hybridization between the $x$ and $y$ layers vanishes at the Fermi level, it remains finite at lower energies. In Fig. (3B), the hybridization increases as inter-layer hopping increases. Furthermore the Mott band approaches the Fermi-energy in Fig. (3B) Finally, when we transition to a layer uniform phase, we lose the layer character of the electrons as shown in Fig. (3C). We present the upper Hubbard bands as an inset of each panel, and they are separated from the lower Hubbard band by an energy gap of $\sim U$. Interestingly, the upper Hubbard bands remain entirely unhybridized in the LSMP phase.
@@
Figure 4: Layer resolved spectral function of the lower Hubbard bands for $\lambda=0.06t$ and $n=1.72$, corresponding to the blue circle in the phase diagram in Fig. (1A) displays the intra-layer spectral function $A_{\alpha\alpha}$ as a function of energy $\omega-\mu$ for different high symmetry $k$-points. The spectral for each $k$-points are shifted by $1.5t$ for clarity. (B) Shows the inter-layer spectral function $A_{\alpha\beta}$ for the same parameter. Here the results for each $k$-points are shifted by $0.3t$ for clarity.
@@
Therefore, we reveal the characteristics of the bands in the LSMP and LUP phases. First, in the LSMP phase, one of the bands flattens and goes below the Fermi level. Secondly, in the LSMP phase, the two bands have weak hybridization and retain their layer character. Thirdly, the hybridization occurs only around specific momentum points where the two bands overlap. However, such hybridization remains appreciable for any finite $\lambda$, indicating the absence of a complete breakdown of hybridization in the LSMP phase.
@@ III.3 Fermi surface
Figure 5: (A) Displays the spectral function $A(\mathbf{k},E_{F})$ at $\lambda=0.06t$ in the LSMP phase for $n=1.72$ (blue circle on Fig. (1A) ). (B) Depicts the same for $\lambda=0.2t$ (blue square on Fig. ( 1A)), in the metastable region of the LSMP. (C) Fermi surface of the layer uniform phase at $\lambda=0.36t$ (blue star on Fig. (1A)). In the LSMP, only a single Fermi sheet is observable, whereas in the metallic phase, two sheets are visible. Furthermore, the Fermi surface area is smaller in the LSMP phase compared to the LUP. (D) Illustrates the evolution of the Fermi volume $\mathcal{A}_{s}$ as a function of inter-layer hopping for different electron densities, showing a jump at the transition. In the LSMP phase, the Mott layer does not contribute to the Fermi volume since one band remains filled.
In Fig. (5A), we illustrate the Fermi surface contour, for $\lambda=0.06t$ at $n=1.72$ in the LSMP phase. In the LSMP phase, a single Fermi sheet is evident since the Mott layer remains fully occupied. With increasing inter-layer hopping there is little change in the Fermi surface as shown in Fig. (5B). In contrast, both bands contribute to the Fermi surface in the metallic regime at $\lambda=0.36t$. Additionally, there is a significant increase in the size of the Fermi surface from the LSMP to the uniform phase.
Moreover, we monitor the increase in the volume of the Fermi surface across the transition with inter-layer hopping for various total electron densities in Fig. (5D). The volume exhibits a distinct jump at $\lambda_{c}$, which is more pronounced for lower $n$. From the evolution of the Fermi surface, the LSMP phase appears to be in a Kondo breakdown-like phase. However, we have shown in the previous section that hybridization of the two layers remains non-zero in the LSMP phase.
A characteristic feature of non-Fermi liquid behavior is the violation of the Luttinger theorem, which asserts that the volume of the Fermi surface must equal the electron density. This theorem [60] applies to any system characterized by a low-energy theory of a Fermi liquid. As the composite excitations in our system are non-fermionic, a violation of the Luttinger theorem is expected [61, 62]. Consequently, our calculations demonstrate that the Luttinger theorem is violated in both the LSMP and strongly correlated LUP phase.
Hence, we observe a small Fermi surface in the LSMP to a large Fermi surface in the LUP phase. As a result, we detect a violation of the Luttinger theorem in the layer selective phase. Despite the sharp increase in the Fermi volume in the LUP phase, the Luttinger theorem continues to be violated. Thus, the traditional Fermi liquid picture is not applicable in the strongly correlated Hubbard model within this formalism.
@@ III.4 Density of states
Figure 6: (A) Presents the low-energy density of states at $n=1.72$ for several $\lambda$ ranging from 00 to $0.42t$ with a $0.02t$ step. Each $\lambda$ curves are shifted vertically by $0.05t$ for clarity.$\lambda<0.25t$ corresponds to the LSMP phase. (B), shows the layer resolved density of states at $\lambda=0.16t$. Here, the $y$ layer approaches the Mott character, indicated by the reduced DOS at Fermi energy, whereas the $x$ layer remains metallic. The peak at the negative energy comes from the Mott layer. A progressive flattening of such peak of the LSMP occurs in (A) with increasing inter-layer hopping. (C) Shows the value of DOS at Fermi-energy, which manifest a jump around $\lambda=0.24t$, indicating the transition from LSMP to the uniform phase. (D) Tracks the energy associated with the Mott peak $\omega_{\rm max}$ (black dots in A).
We now focus on the density of states (DOS). In Fig. (6A), we illustrate the low-energy density of states for $n=1.72$ across different $\lambda$ values. The curves for different $\lambda$ values are shifted vertically by $0.05t$ for clarity. Initially, at weak $\lambda$, the DOS features a prominent peak at negative energies (shown in black dots), with a finite DOS at the Fermi energy $E_{F}$. The presence of mobile electrons in one layer maintains an ungapped DOS at $E_{F}$. As inter-layer hopping increases, so does the DOS at Fermi energy. The transition from the LSMP to the LUP state is marked by a jump in the evolution of $D(E_{F})$, as shown in Fig. (6C).
In Fig. (6B), we depict the layer-resolved DOS for ${\lambda=0.17}$. The Mott layer $y$ displays a distinct peak at negative energy, accompanied by a small DOS at $E_{F}$. This peak signifies the energy required to extract electrons from the Mott layer. We track the evolution of the Mott peak $\omega_{\rm max}$ as a function of inter-layer hopping in Fig. (6D). The energy $\omega_{\rm max}$ shows a smooth reduction. As a result of the hybridization between heavy Mott electrons and itinerant electrons, the peak flattens, making it challenging to discern the layer character of electrons prior to the transition to LUP.
Therefore, the LSMP phase features a reduced DOS at Fermi energy. The transition from the LSMP to the uniform phase is marked by a sharp increase in the states at the Fermi energy at a critical inter-layer hopping. Additionally, the LSMP phase displays a strong negative bias peak, which flattens, indicating the hybridization of Mott electrons with itinerant electrons. Notably, the flattening of the Mott peak with increasing inter-layer hopping occurs more smoothly than the sharp transitions observed in other quantities during the phase transition. This indicates that hybridization among the layers ultimately eliminates the layer differentiation.
@@ III.5 Quasiparticle residue
Figure 7: (A)) Quasiparticle weight for several density as a function of $\lambda$, the errorbars are standard deviation when averaging over Fermi momentum. We overlapped with points the value of the jumps observable in $N_{k}$. (B) layer resolved quasiparticle weight. The Mott layer has a vanishing quasiparticle residue for finite $\lambda$.
@@
We define the average electronic occupancy at a given momentum $k$, denoted as $N_{k}$, as follows:
$N_{k}=\dfrac{1}{4}\sum\limits_{\alpha\sigma}\langle c^{\dagger}_{\alpha\sigma}c_{\alpha\sigma}\rangle$ (33)
We report the typical shape of $N_{k}$ along the $\Gamma$ to $M$ high symmetry line in Fig. (9). By definition of the Green’s function in Eq. 80, when crossing the Fermi contour a discontinuity of the order of $Z$ is observed in momentum space. One drop is observed for the LSMP, and the momentum resolved density does not vanish at $M$ due to the presence of the fully filled Mott band. Two drops are seen in the LUP due to the two Fermi contours of the phase. In this case, $Z$ is defined as the sum of these two drops. We report the $Z$ measured this way as a function of $\lambda$ on Fig (7A) as dots. There is a good agreement between the two ways of computing $Z$. Since there is no Fermi surface at half-filling, and that $Z$ is a quantity defined at the Fermi energy, $Z=0$ at half-filling. As electron density $n$ is decreasing, $Z$ increases. Therefore, we can also interpret the quasi-particle weight as a measurement of how close the system is from Mott regime, where interaction affect the system the most.
Fig. (7B) displays the layer resolved quasi-particle weight. In the LSMP the Mott layer has an almost (exactly at $\lambda=0$) vanishing quasiparticle weight due to the localized nature of the electron. Since the other layer is further away from half-filling than the average doping its associated $Z$ increases from the average value.
@@ III.6 Phase diagram with Hubbard interactions
Figure 8: Here, we present the phase diagram as a function of intra-layer interaction strength $U$ and $\lambda$ for $n=1.85$. The colored area shows the region where the LSMP phase is energetically favorable, whereas the LUP phase is favorable in the white region. The LSMP is a metastable state between the thick and dashed lines.
@@
Finally, In Fig. (8), we present the phase diagram when the on-site repulsion $U$ is varied for electron density $n=1.85$. The colored region indicates where the LSMP phase is more favorable than the uniform phase. On the other hand, the LUP phase is stable in the white areas. The LSMP is a metastable state between the thick and dashed lines.
@@
For lower $U$, there is a weak increment of critical inter-layer hopping $\lambda_{c}$ where the LSMP phase becomes unfavorable. For large $U>15$ the same $\lambda_{c}$ reduces with increasing $U$. Thus, the critical $\lambda_{c}$ shows a nonmonotonic behavior with Hubbard interaction $U$.
@@
To qualitatively understand this result, we examine the destabilization of the LSMP phase. Initially, we assume double occupancy is forbidden due to strong repulsion. First, we focus on comprehending the decrease of $\lambda_{c}$ with increasing Hubbard interaction.
@@
In the LSMP phase, all the Mott layer sites are singly occupied, whereas the metallic layer has itinerant electrons and vacancies. When $\lambda\neq 0$, electrons can hop from the Mott to the metallic layers if the metallic site is vacant. A stable LSMP phase depends on refilling the holes in the Mott layer without significant hybridization. Without recombination, holes will proliferate in the Mott layer, leading to the breakdown of the layer differentiation. The hole created in the Mott layer is heavy and can be assumed to be approximately immobile.
@@
Two scenarios are possible once an electron hops to the metallic layer: it either immediately hops back to the Mott layer or transfers to a neighboring site within the same layer. The stability of the LSMP phase depends on the ratio between $t$ and $\lambda$ for a fixed $U$ and electron density.
@@
In the $\lambda\ll t$ regime, electrons are more likely to hop between neighboring sites of the same layers rather than between layers. Since double occupancy is forbidden, recombination can only happen when any electron on the metallic layer returns to the site at which Mott layer has the hole and goes back to the Mott layer. The mobility of electrons in the metallic layers is crucial in such a recombination pathway, which is relatively enhanced for lower $U$. Thus, the LSMP phase survives for a broader range of $\lambda$ when $U$ is lower.
@@
In the $\lambda\gg t$ regime, inter-layer hopping between the $x$ and $y$ layers dominates. This leads to strong hybridization between the $x$ and $y$ layers, which destroys the LSMP phase.
@@
Finally, we focus on understanding why $\lambda_{c}$ decreases for $U<15$. In this regime, double occupancy contributes to destabilizing the LSMP ordering. As $U$ decreases, double occupancies become more likely, increasing hybridization between the Mott and itinerant layers. Consequently, a metastable LSMP phase can still be present at lower $U$ due to the competing processes that are favored when lowering $U$. However, the LSMP becomes energetically less favorable compared to the uniform phase.